# Detection and Tracking of MAVs Using a Rosette Scanning Pattern LiDAR

### v2

@@ Abstract
The use of commercial Micro Aerial Vehicles (MAVs) has surged in the past decade, offering societal benefits but also raising risks such as airspace violations and privacy concerns. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority. In this study, we tackle this challenge, by using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor. The presented method utilizes a particle filter with a velocity component for the detection and tracking of the drone, which offers added re-detection capability. A Pan-Tilt platform is utilized to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense. The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments. Our approach achieved accuracy on par with the state-of-the-art indoor method while increasing the maximum detection range by approximately 80% beyond the state-of-the-art outdoor method.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I INTRODUCTION
@@
The state-of-the-art drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24]. Dogru et al.[13] achieve a detection distance of $70\;m$ outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/. This detection distance can only be extended to $100\;m$ by adding a retroreflector to the drone. Catalano et al.[24] use a LiVOX Horizon [^5] https://www.livoxtech.com/ to achieve a few $cm$ accuracy using a method based on Kalman Filter which has no re-detection capability in case of lost tracking. Their method was only tested indoors so there is no information on the maximum detection distance.
Our proposed algorithm uses a particle filter with a velocity component to achieve fast detection and tracking with added re-detection capability. Leveraging the point distribution of the LiVOX AVIA LiDAR [^6] https://www.livoxtech.com/avia, the proposed method keeps the MAV in the middle of the sensor’s FoV, where the measurements are most dense, using a pan-tilt platform. With this approach we achieved accuracy on par with the state-of-the-art indoor method using a similar sensor [24], while increasing the maximum detection range approximately $80\%$ beyond the state-of-the-art outdoor method [13] by detecting drones up to $120m$ away without attached retroreflector. In addition, the number of points reflected from the drone at the same distance increased several folds over the full range.
@@
To summarize, this paper enhances the state-of-the-art in LiDAR-based drone tracking with the following contributions:
@@
  • The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments [^7] Video demonstration of experiments at https://youtu.be/VdePm3on8vo.
The remainder of the paper is organized as follows: related work is presented in section II, the proposed algorithm in section III, the experiments in sections IV and V, and finally the conclusion in section VI.
II RELATED WORK
@@
In [36] a method is proposed for finding MAVs in a point cloud generated by a rosette scanning LiDAR on a moving robot. This approach is further improved by adding an Kalman Filter based tracking method and extensive evaluation in [24]. In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about $30\;m$. However, this method does not allow for regaining tracking once lost. Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor, thereby maximizing the number of returned points. Additionally, our approach incorporates a robust tracking method capable of regaining tracking.
[…]
LiDAR sensor performance in outdoor environments is highly influenced by weather conditions [37], [38]. According to [39] the detection range of LiDAR sensors is limited to $25\;m$ in foggy conditions when the meteorological visibility is under $40\;m$. Based on this a significant reduction in the detection distance and the number of detected points can be expected for outdoor tests under foggy weather conditions. While our experiments confirmed this trend, our method could still track the drone to a little more than $50\;m$.
III THE PROPOSED ALGORITHM
The flowchart of our method is presented in Fig. 1, where the main steps of the algorithm are shown. Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section III-A). This is used to filter the static background from the point cloud to detect the target object.
@@ III THE PROPOSED ALGORITHM
Next, as part of the object tracking, the particle filter[40] performs the prediction step. If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section III-B). The three steps of the particle filter (predict, update, and resample) are explained in detail in Section III-C.
Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section III-D). As a result, the measured point density reflected from the drone is maximized.
III-A Building the Background Model
In our application, the background is considered static throughout the drone tracking, hence it is built once at the beginning of the algorithm. Consequently, the proposed method considers everything that is not recorded in the background a drone. The turret is set to scanning mode and performs movements to scan a pre-defined area. Because of the LiDAR’s non-repetitive way of scanning, the point density of the whole scene will increase the longer the scanning time is.
@@ III-A Building the Background Model
Based on the position and orientation of the LiDAR, the incoming point cloud is transformed into the static coordinate system of the background model. The points are filtered based on position thresholds; the ground plane, near points, and far away points are removed. After filtering, the point cloud is inserted into the OctoMap of the background model. As the last step of building the background model, after inserting all the collected point clouds, inflation is performed to account for measurement noise (e.g. vibrations of the turret base) as well as small changes in background objects (such as leaves moving on the trees).
III-B Processing New Measurements
The point measurements are received in the Cartesian frame of the LiDAR and transformed into the World frame based on the turret’s pan and tilt angles. Then incoming point cloud is filtered as during the background building: the ground plane, near points, and far away points are removed. And finally, the background model is subtracted. There are two additional filters to remove the reflected points that are not close enough to each other and cannot belong to the target object. The filters are the RadiusOutlierRemoval and StatisticalOutlierRemoval algorithms from the Point Cloud Library (PCL) [42].
III-C Object Tracking Using Particle Filter
@@
The particle filter is implemented following Algorithm 2. in [43]. Each particle represents a possible state of the drone, where the state is the 3D location and 3D velocity of the drone ($s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$).
III-C1 Predict
@@
First, the position of each particle is predicted based on its previous position, velocity, and added process noise. The prediction noise $p_{\{x,y,z\},k}$ is sampled from a 00 mean Gaussian distribution with $\sigma_{pred}$ standard deviation and $T$ denotes the period of the particle filter.
$p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pred}^{2})$ (1)
$\begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+p_{z,k}\end{bmatrix}$ (2)
@@
The velocities are computed only if the drone is tracked. The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pred}$. The $\sigma_{particles}$ value increases with $\sigma_{pred}$ plus the additional effect of the velocity in every iteration when the drone is lost. The drone is considered tracked if it is under $\sigma_{threshold}=2\cdot\sigma_{pred}$. In this case, the velocities are predicted, or else they are set to 00.
$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}(x_{k}-x_{k-1})/T\\ (y_{k}-y_{k-1})/T\\ (z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
III-C2 Update
@@
The weight of each particle is updated by the product of the previous weight and the Gaussian distribution with $\sigma_{m}$ standard deviation and 00 mean evaluated at $d_{i}$, where $\sigma_{m}$ represents the measurement noise and $d_{i}$ is the distance from the $i$th particle to the closest LiDAR measurement point. This product is normalized, so the sum of $w_{i}$ will add up to $1$, where $n$ is the number of particles.
$w_{i,k}=\frac{w_{i,k-1}\cdot\mathcal{N}(d_{i};0,\sigma_{m}^{2})}{\sum_{i}^{n}w_{i}}$ (4)
III-C3 Resample
@@
III-C4 Particle initialization and reset
The particles are initialized randomly in a cylinder, whose parameters (3D location, radius, and height) are set based on the detection area. The particles can be reset to a new cylinder during runtime, thus giving the operator the capability to reset the tracking to another object. During tracking $\sigma_{particles}<\sigma_{threshold}$ which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$ which is almost certainly a collision of the objects.
III-D Turret control
@@
III-E Implementation details

### v3

@@ Abstract
The use of commercial Micro Aerial Vehicles (MAVs) has surged in the past decade, offering societal benefits but also raising risks such as airspace violations and privacy concerns. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority. In this study, we tackle this challenge, by using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor. The presented method utilizes a particle filter with a velocity component for the detection and tracking of the drone, which offers added re-detection capability. A pan-tilt platform is utilized to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense. The system’s tracking capabilities (both in coverage and distance), as well as its accuracy are validated and compared to State Of The Art (SOTA) models, demonstrating improved performance, particularly in terms of coverage and maximum tracking distance. Our approach achieved accuracy on par with the SOTA indoor method while increasing the maximum detection range by approximately $85\;\%$ beyond the SOTA outdoor method to 130m130\;m. Additionally, our method yields at least a twofold increase in track coverage and returned point counts.
Index Terms:
Drones, Tracking, Trajectory tracking, Particle filters, Sensors
This work has been published in IEEE ACCES https://doi.org/10.1109/ACCESS.2025.3596857
@@
The State Of The Art (SOTA) drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24]. Dogru et al.[13] achieve a detection distance of 70m70\;m outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/. This detection distance can only be extended to 100m100\;m by adding a retroreflector to the drone. Catalano et al.[24] use a LIVOX Horizon [^5] https://www.livoxtech.com/ to achieve a few $cm$ accuracy using a method based on Kalman Filter which has no re-detection capability in case of lost tracking. Their method was only tested indoors so there is no information on the maximum detection distance.
Our proposed algorithm uses a Particle Filter (PF) with a velocity component to achieve fast detection and tracking with added re-detection capability. Leveraging the point distribution of the LIVOX AVIA LiDAR [^6] https://www.livoxtech.com/avia, the proposed method keeps the MAV in the middle of the sensor’s FoV, where the measurements are most dense, using a pan-tilt platform. With this approach we achieved accuracy on par with the SOTA indoor method using a similar sensor [24], while increasing the maximum detection range approximately $85\%$ beyond the SOTA outdoor method [13] by detecting drones up to 130m130\;m away without attached retroreflector. Additionally, our method yields significantly higher track coverage, meaning the drone is tracked for a greater percentage of its trajectory. Furthermore, the number of points reflected from the drone at the same distance increased several folds over the full range. We provide detailed experiments and comparisons to support these claims and we also analyze how much each component of our algorithm contributes to the observed improvement in tracking distance.
@@
To summarize, this paper enhances the SOTA in LiDAR-based drone tracking with the following contributions:
@@
  • The system’s tracking capabilities (both in coverage and distance), as well as its accuracy are validated and compared to SOTA models [13, 24], demonstrating improved performance, particularly in terms of coverage and maximum tracking distance. A video is shared of the experiments. [^7] Video demonstration of experiments at https://youtu.be/ghfjqAnDuag
  • We release the source code of the proposed algorithm and the datasets used in the experiments. [^8] Source code at https://github.com/gsanya/LIVOX_MAV_track
The remainder of the paper is organized as follows: related work is presented in Section I, the proposed algorithm in Section II, the experiments in Section III, and finally the conclusion in Section IV.
I RELATED WORK
@@
Other than detection, recent works have also focused on visual tracking. The rapid development of deep learning-based visual tracking methods is explored in [28]. Learned motion models have been shown to outperform traditional Kalman Filter-based approaches, particularly in handling non-linear motion [29]. However, these methods typically rely on dense visual input and are not well-suited for sparse point clouds with a limited number of measurement points. Moreover, these works generally target generic object tracking rather than specifically addressing the challenges of tracking small, fast-moving drones.
@@
In [38] a method is proposed for finding MAVs in a point cloud generated by a rosette scanning LiDAR on a moving robot. This approach is further improved by adding a Kalman Filter based tracking method and extensive evaluation in [24]. In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about 30m30\;m. However, this method does not allow for regaining tracking once lost. Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor FoV, thereby maximizing the number of returned points. Additionally, our approach incorporates a robust tracking method capable of regaining tracking.
In the case of a fixed LiDAR configuration, the usability range in [12] is less than 10m10\;m for the VLP16 sensor and 3550m35-50\;m for the HDL-64 sensor. In [13], it was demonstrated that few measurements are sufficient to track the targets and the maximum detection range was significantly increased with their method. Their approach achieved a maximum detection distance of 70m70\;m with sparse point clouds, and by using retroreflectors, this was extended to 100m100\;m. The applied retroreflectors did not increase the size of the drone but increased the number of reflected points from the drone. Our method greatly outperforms this baseline solution in similar weather conditions with a maximum tracking distance of more than 130m130\;m without retroreflector.
LiDAR sensor performance in outdoor environments is highly influenced by weather conditions [39], [40]. According to [41] the detection range of LiDAR sensors is limited to 25m25\;m in foggy conditions when the meteorological visibility is under 40m40\;m. Based on this a significant reduction in the detection distance and the number of detected points can be expected for outdoor tests under foggy weather conditions. While our experiments confirmed this trend, our method could still track the drone to a little more than 50m50\;m. These findings reflect incidental observations rather than a systematic study of weather effects.
II THE PROPOSED ALGORITHM
The flowchart of our method is presented in FIGURE 1, where the main steps of the algorithm are shown. Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section II-A). This is used to filter the static background from the point cloud to detect the target object.
@@ II THE PROPOSED ALGORITHM
Next, as part of the object tracking, the particle filter[42] performs the prediction step. If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section II-B). The three steps of the particle filter (predict, update, and resample) are explained in detail in Section II-C.
Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section II-D). As a result, the measured point density reflected from the drone is maximized.
II-A Building the Background Model
In our application, the background is considered static throughout the drone tracking, hence it is built once at the beginning of the algorithm. Consequently, dynamic objects such as cars, moving foliage, and pedestrians can lead to false positives. To address this, we apply simple mitigation strategies that have proven effective in practice. While we also explored a transformer-based background segmentation approach in [43], it is not used in this work and falls outside the scope of this paper. The turret is set to scanning mode and performs movements to scan a pre-defined area. Because of the LiDAR’s non-repetitive way of scanning, the point density of the whole scene will increase the longer the scanning time is.
@@ II-A Building the Background Model
Based on the position and orientation of the LiDAR, the incoming point cloud is transformed into the static coordinate system of the background model. The points are filtered based on an operator set bounding box, which is set to remove very far away points, the ground, and ground adjacent points, such as pedestrians and cars. After filtering, the point cloud is inserted into the OctoMap of the background model. As the last step of building the background model, after inserting all the collected point clouds, inflation of the background model is performed to account for measurement noise (e.g. vibrations of the turret base) as well as small changes in background objects (such as moving foliage).
The primary objective of the background removal step is to eliminate the majority of points from the LiDAR measurements that do not correspond to the target. Although some false positives may persist, the tracking algorithm has sufficient robustness to handle these residual points.
II-B Processing New Measurements
The point measurements are received in the Cartesian frame of the LiDAR and transformed into the World frame based on the turret’s pan and tilt angles. The incoming point cloud is filtered during the background building: the ground plane, near points, and far away points are removed based on the same bounding box used at background model building. And finally, the background model is subtracted. There are two additional filters to remove the reflected points that are not close enough to each other and cannot belong to the target object. The filters are the RadiusOutlierRemoval and StatisticalOutlierRemoval algorithms from the Point Cloud Library (PCL) [45].
II-C Object Tracking Using Particle Filter
@@
The particle filter is implemented following Algorithm 2. in [46]. Each particle represents a possible state of the MAV. Three different state representations and transition models were implemented and tested, namely: a delta position based estimator, one using a constant velocity motion model and one using a constant acceleration motion model. The state vector $s$ varies depending on the chosen model. For the delta position and constant velocity models $s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$ while for the constant acceleration model $s=[x,y,z,v_{x},v_{y},v_{z},a_{x},a_{y},a_{z}]^{T}$. The position estimation of the drone is calculated by averaging the particle positions. In the upcoming sections we present the equations governing the predict, update, and resample steps of the particle filter for each of the three models.
II-C1 Predict
@@
First, the position of each particle is predicted based on its previous position, velocity, acceleration, and added process noise as in (1). The position prediction noise $p_{\{x,y,z\},k}$ is sampled from a 0 mean Gaussian distribution with $\sigma_{pos}$ standard deviation, $T$ denotes the period of the particle filter, and $a_{\{x,y,z\},k-1}=0$ in case of the delta position and constant velocity models.
$\begin{gathered}p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pos}^{2})\\ \begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+0.5\cdot a_{x,k-1}\cdot T^{2}+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+0.5\cdot a_{y,k-1}\cdot T^{2}+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+0.5\cdot a_{z,k-1}\cdot T^{2}+p_{z,k}\end{bmatrix}\end{gathered}$ (1)
@@
The velocities are computed in two different ways. In case of the delta position model, the term is dependent on the status of the tracking. The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pos}$. The $\sigma_{particles}$ value increases with $\sigma_{pos}$ plus the additional effect of the velocity in every iteration when the drone is lost.
The MAV is considered to be successfully tracked if $\sigma_{particles}$ remains below a threshold defined as $\sigma_{threshold}=1.5\cdot\max(\sigma_{pos},\sigma_{m})$, where $\sigma_{m}$ is the measurement uncertainty. This condition ensures that when the measurement is less reliable than the position estimate, the resulting estimate reflects the greater uncertainty. We scale the velocity with the scaling factor $\rho$ that is inversely proportional to that uncertainty.
$\rho=\begin{cases}1,&\text{if }\sigma_{particles}<\sigma_{threshold}\\ \frac{\sigma_{pos}}{\sigma_{particles}},&\text{otherwise}\end{cases}$ (2)
$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}\rho\cdot(x_{k}-x_{k-1})/T\\ \rho\cdot(y_{k}-y_{k-1})/T\\ \rho\cdot(z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
In case of the constant velocity and constant acceleration models, the velocity is computed as in (4-5), where $a_{\{x,y,z\},k-1}=0$ in case of the constant velocity model.
$p_{\{v_{x},v_{y},v_{z}\},k}\sim\mathcal{N}(0,\sigma_{vel}^{2})$ (4)
$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}v_{x,k-1}+a_{x,k-1}\cdot T+p_{v_{x},k}\\ v_{y,k-1}+a_{y,k-1}\cdot T+p_{v_{y},k}\\ v_{z,k-1}+a_{z,k-1}\cdot T+p_{v_{z},k}\end{bmatrix}$ (5)
Finally, in case of the constant accelartion model the acceleration is updated as in (6-7).
$p_{\{a_{x},a_{y},a_{z}\},k}\sim\mathcal{N}(0,\sigma_{acc}^{2})$ (6)
$\begin{bmatrix}a_{x,k}\\ a_{y,k}\\ a_{z,k}\end{bmatrix}=\begin{bmatrix}a_{x,k-1}+p_{a_{x},k}\\ a_{y,k-1}+p_{a_{y},k}\\ a_{z,k-1}+p_{a_{z},k}\end{bmatrix}$ (7)
II-C2 Update
@@
The weight of each particle is updated when a new measurement arrives based on (8). The new weight is calculated using a Gaussian distribution with $\sigma_{m}$ standard deviation, which represents the uncertainty of the lidar sensor measurement. This distribution is evaluated at $d_{i}$, which is the distance between the $i$th particle and the closest LiDAR measurement point. The weights are then normalized (9) so that the sum of $\tilde{w_{i}}$ will add up to 1, where $n$ is the number of particles.
$w_{i}=\mathcal{N}(d_{i};0,\sigma_{m}^{2})$ (8)
$\tilde{w_{i}}=\frac{w_{i}}{\sum_{j}^{n}w_{j}}$ (9)
II-C3 Resample
@@
II-C4 Particle initialization and reset
The particles are initialized randomly in a cylinder, whose parameters (3D location, radius, and height) are set based on the detection area. The particles can be reset to a new cylinder during runtime, thus giving the operator the capability to reset the tracking to another object. During tracking $\sigma_{particles}<\sigma_{threshold}$, which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$, which is almost certainly a collision of the objects.
II-D Turret control
@@
The pan ($\alpha$) and tilt ($\beta$) goal angles of the turret are calculated based on the current drone position estimate as in (10).
$\alpha$ $=\tan^{-1}\left(\frac{y_{mav}}{x_{mav}}\right)$ (10)
$\beta$ $=-\tan^{-1}\left(\frac{z_{mav}}{\sqrt{(x_{mav})^{2}+(y_{mav})^{2}}}\right)$
II-E Implementation details